Calculate the longest and shortest wavelengths (in m) of light emitted by electrons in the hydrogen atom that begin in the

n = 3

state and then fall to states with smaller values of n.

longest wavelength m

shortest wavelength m
Answer #1

Shorter wavelength = 1.02518× 10-7 this is for transition from n=3 to n=1

Larger wavelength = 6.561142×10-7 this is for transition from n=3 to n=2